I see, I need to teach you some math guys. Here we go.
Let's calculate the probability of failure when you craft 4 pieces 1 by 1, like when all of them are burned. It's 0.78 * 0.78 * 0.78 * 0.78 = 0.37.
The probability of a failed craft when you craft 4 at once is 0.22.
Now you see it's (0.37 - 0.22) * 100% = 15% difference. It's quite a lot. 🐶 Post #610
22
Green Forest 🌲 How would you craft gifts?
- 👍 2